# A Comprehensive Evaluation of Augmentations for Robust OOD Self-Supervised Contrastive Phonocardiogram Representation Learning

### 18 Mar 2024 v2

@@ 3.2 Self-Supervised Contrastive Learning for PCG Representation Learning
For the contrastive prediction task, we employ the contrastive NT-Xent loss function introduced in [27] and implemented in the SimCLR framework [67]. Specifically, given a set $\{\tilde{x}_{k}\}$ which includes an augmented positive pair $\tilde{x}_{i}$ and $\tilde{x}_{j}$, a model trained via contrastive SSL aims to identify $\tilde{x}_{j}$ in $\{\tilde{x}_{k}\}_{k}\neq i$ for a given $\tilde{x}_{i}$. The contrastive loss for a positive pair $(\tilde{x}_{i},\tilde{x}_{j})$ is formulated as follows:
@@ 3.4 Evaluation Protocol
In Step 4, a classification head with learnable weights is attached to the pretrained encoder which is then trained on data from one of the labeled datasets (Pascal, PhysioNet2016 or PhysioNet2022). Specifically, each of these datasets is split into training, validation and test splits with a standard 70%-20%-10% ratio and the trained model is evaluated on the test split only. Since each model is evaluated on test data originating from the same distribution as its training data, we call this the ”In-Distribution” evaluation. Inspired by recent work in biosignal domain generalization [9], we also design an evaluation protocol for assessing the robustness of a trained classifier on out-of-distribution data, which are drawn from entirely different datasets. Since the data has been completely left-out during downstream training, the above process is followed for evaluating the generalization ability of a model. Specifically, we now also freeze the weights of the attached classification layers and test the model on data drawn from the remaining 2 labeled datasets. For example, if the network is trained for the Pascal downstream task, it is evaluated on both in-distribution data (Pascal) and on out-of-distribution data (PhysioNet2016 and PhysioNet2022). We call this step the out-of-distribution or OOD evaluation step. However, in order to evaluate the same model on each separate dataset, all data must share common labels. To that end, we implement two different downstream task types: ‘All’ and ‘Binary‘. In the ‘All‘ downstream task, the original labels of each dataset are assigned to each audio window, while in the ‘Binary’ task a signal is labeled as either ‘normal’ or ‘abnormal’, as described in Section 3.1. In the case of the PhysioNet2016 dataset, the two tasks match as it introduces a binary classification problem in the first place. For a fair comparison, all test data splits remain the same across each evaluation scheme. Finally, to assess the pretrained backbone encoder (and therefore the applied augmentations) on all available data, we repeat Steps 4 and 5 with the remaining datasets in a round-robin fashion (step 6 of the proposed framework).
@@ 3.5 Implementation Details
In this section, we describe all the implementation details regarding the execution of our experiments. First of all, for the backbone encoder model we implement the same model used in [24] and [68]. The encoder model is a 5 layer convolutional neural network, the architecture of which is depicted in Figure 2 (left).
For the contrastive SSL pretraining phase, we train the conv layers of the network by projecting the output of the last layer into a 128-D space via a dense or fully-connected layer and minimize the NT-Xent loss (eq. 1). For the cosine similarity (eq. 2) we use a temperature of 0.1, as in [24], while the initial batch size is set to 256 (leading to 512 after augmentation occurs). The SSL training epochs are set to a maximum of 200 epochs, based on an early stopping set on the validation split with a patience of 10 epochs. Additionally, to accommodate the large batch size we employ the LARS optimizer [69] with a linear learning rate warm-up of 20 epochs, up to 0.1, followed by a cosine decay of 1%.
During the downstream task, the dense projection layer of the pretrained encoder is discarded and all convolutional weights are frozen. Next, a 3-layered classification head is appended to the pretrained network, as shown in Figure 2 (right). Dropout layers are also included between the dense layers of the classification head for further improving model robustness. The output size of the last decision layer depends on the downstream task type and the label size in each dataset. For the ‘Binary’ type and the PhysioNet2016 dataset, the output size is set to 1. For the ‘All’ type, the output size is set to 5 or 3, for the Pascal and PhysioNet2022 datasets respectively. In all cases, the classification head is trained via the appropriate cross-entropy loss and the Adam optimizer. Additionally, the learning rate is set to $10^{-4}$ and the batch size is set to 32. Once again, we train the model for 100 epochs at most, using early-stopping with 20 epochs patience. The same CNN model is used for the fully-supervised baseline. All experiments are implemented in Tensorflow 2.12 [70] and trained on a SLURM [71] cluster, containing 4$\times$40GB NVIDIA A100 GPU cards, split into 8 20GB virtual MIG devices.
@@ 4.1 Augmentation and Transformation Evaluation
Table 2: This table presents the experimental results of the proposed evaluation protocol, as specified in section 3.4 and Figure 1. Top results are specified in bold face. The first column of the table indicates the datasets on which the SSL backbone was pretrained on. As in Fig. 3.4, E, F, Pa, P16 and P22, indicate data from the Ephnogram, Fpcgdb, Pascal, Physionet16 and Physionet22 datasets respectively. The SSL and fully-supervised baseline models are trained and evaluated on In-Distribution data from the respective dataset. The presented models are the best-performing models, based on results from the In-Distribution test data split. Once trained, they are then evaluated on out-of-distribution data originating from unseen datasets. The ‘All’ labels indicate that the PCG signals are classified based on the dataset’s original labels (P16 has only binary labels), whereas ‘Binary’ indicates that the signals are classified between normal and abnormal. The SSL-NoDs model refers to the backbone network which has not ‘seen’ data from the downstream task dataset during pretraining. The adopted augmentation combinations for each model are reported in Section 4.1.
@@
While the above augmentations agree, for the most part, with the distributions and number of occurrences in Fig. 3, they also support the findings in several previous contrastive SSL works [63, 72], in which stronger augmentations lead to improved results.
To further validate our results and to draw concrete conclusions regarding the importance of each applied transformation, we calculate Cohen’s $d$ [73] for each implemented function. Cohen’s $d$ provides a standardized measure of the effect size of each augmentation and allows the comparison of their magnitude. Given two groups $x_{1}$ and $x_{2}$, Cohen’s $d$ is the difference in their means divided by the pooled standard deviation ($s$). Specifically, Cohen’s $d$ is defined as:
@@ 4.3 Discussion
By diving further into the results, the true benefits of the proposed method begin to show. First of all, the drop in effectiveness between in-distribution and OOD data is apparent. Additionally, in all cases the baseline models appears to have overfit their training data distribution and fails to maintain its classification ability across distinct datasets. On the other hand, the models which have been pretrained via SSL yield far superior results. In all cases, the models trained with the proposed framework surpass the baseline by 11.68% on average and even outperform it by 21.04% in the case of the PhysioNet22 OOD task. Even more so, the SSL models were surprisingly able to surpass both the effectiveness of the baseline and their ‘In-Distribution’ counterpart, in the case of the Pascal OOD evaluation. Specifically, the downstream models trained on data from PhysioNet2016 and PhysioNet2022 were able to yield better results than the fully-supervised model trained solely on Pascal data. Furthermore, even though there is a decrease in effectiveness when the OOD datasets are left out during model pretraining, the SSL models continue to outshine the baseline across the board, with the exception of the case where the downstream task is the PhysioNet2022 data and models are evaluated on OOD data from Pascal. Finally, we would like to note that the above findings are aligned with the results of our preliminary SSL method [24] submitted in the PhysioNet2022 challenge, where the proposed model was able to avoid overfitting on the available training distribution and demonstrated adequate generalization effectiveness on completely hidden data.

### 05 Apr 2024 v3

@@ 3.2 Self-Supervised Contrastive Learning for PCG Representation Learning
For the contrastive prediction task, we employ the contrastive NT-Xent loss function introduced in oord_representation_2019 and implemented in the SimCLR framework chen2020simple . Specifically, given a set $\{\tilde{x}_{k}\}$ which includes an augmented positive pair $\tilde{x}_{i}$ and $\tilde{x}_{j}$, a model trained via contrastive SSL aims to identify $\tilde{x}_{j}$ in $\{\tilde{x}_{k}\}_{k}\neq i$ for a given $\tilde{x}_{i}$. The contrastive loss for a positive pair $(\tilde{x}_{i},\tilde{x}_{j})$ is formulated as follows:
@@ 3.4 Evaluation Protocol
In Step 4, a classification head with learnable weights is attached to the pretrained encoder which is then trained on data from one of the labeled datasets (Pascal, PhysioNet2016 or PhysioNet2022). Specifically, each of these datasets is split into training, validation and test splits with a standard 70%-20%-10% ratio and the trained model is evaluated on the test split only. Since each model is evaluated on test data originating from the same distribution as its training data, we call this the ”In-Distribution” evaluation. Inspired by recent work in biosignal domain generalization 10233054 , we also design an evaluation protocol for assessing the robustness of a trained classifier on out-of-distribution data, which are drawn from entirely different datasets. Since the data has been completely left-out during downstream training, the above process is followed for evaluating the generalization ability of a model. Specifically, we now also freeze the weights of the attached classification layers and test the model on data drawn from the remaining 2 labeled datasets. For example, if the network is trained for the Pascal downstream task, it is evaluated on both in-distribution data (Pascal) and on out-of-distribution data (PhysioNet2016 and PhysioNet2022). We call this step the out-of-distribution or OOD evaluation step. However, in order to evaluate the same model on each separate dataset, all data must share common labels. To that end, we implement two different downstream task types: ‘All’ and ‘Binary‘. In the ‘All‘ downstream task, the original labels of each dataset are assigned to each audio window, while in the ‘Binary’ task a signal is labeled as either ‘normal’ or ‘abnormal’, as described in Section 3.1. In the case of the PhysioNet2016 dataset, the two tasks match as it introduces a binary classification problem in the first place. For a fair comparison, all test data splits remain the same across each evaluation scheme. Finally, to assess the pretrained backbone encoder (and therefore the applied augmentations) on all available data, we repeat Steps 4 and 5 with the remaining datasets in a round-robin fashion (step 6 of the proposed framework).
@@ 3.5 Implementation Details
In this section, we describe all the implementation details regarding the execution of our experiments. First of all, for the backbone encoder model we implement the same model used in ballas_listen2yourheart_2022 and papapanagiotou_chewing_2017 . The encoder model is a 5 layer convolutional neural network, the architecture of which is depicted in Figure 2 (left).
For the contrastive SSL pretraining phase, we train the conv layers of the network by projecting the output of the last layer into a 128-D space via a dense or fully-connected layer and minimize the NT-Xent loss (eq. 1). For the cosine similarity (eq. 2) we use a temperature of 0.1, as in ballas_listen2yourheart_2022 , while the initial batch size is set to 256 (leading to 512 after augmentation occurs). The SSL training epochs are set to a maximum of 200 epochs, based on an early stopping set on the validation split with a patience of 10 epochs. Additionally, to accommodate the large batch size we employ the LARS optimizer ginsburg2018large with a linear learning rate warm-up of 20 epochs, up to 0.1, followed by a cosine decay of 1%.
During the downstream task, the dense projection layer of the pretrained encoder is discarded and all convolutional weights are frozen. Next, a 3-layered classification head is appended to the pretrained network, as shown in Figure 2 (right). Dropout layers are also included between the dense layers of the classification head for further improving model robustness. The output size of the last decision layer depends on the downstream task type and the label size in each dataset. For the ‘Binary’ type and the PhysioNet2016 dataset, the output size is set to 1. For the ‘All’ type, the output size is set to 5 or 3, for the Pascal and PhysioNet2022 datasets respectively. In all cases, the classification head is trained via the appropriate cross-entropy loss and the Adam optimizer. Additionally, the learning rate is set to $10^{-4}$ and the batch size is set to 32. Once again, we train the model for 100 epochs at most, using early-stopping with 20 epochs patience. The same CNN model is used for the fully-supervised baseline. All experiments are implemented in Tensorflow 2.12 tensorflow2015-whitepaper and trained on a SLURM yoo_slurm_2003 cluster, containing 4$\times$40GB NVIDIA A100 GPU cards, split into 8 20GB virtual MIG devices.
@@ 4.1 Augmentation and Transformation Evaluation
Table 2: This table presents the results of the best performing models (as reported in Section 4.1)Top results are specified in bold face. The first column of the table indicates the datasets on which the SSL backbone was pretrained on. E, F, Pa, P16 and P22, indicate data from the Ephnogram, Fpcgdb, Pascal, PhysioNet16 and PhysioNet22 datasets respectively. The SSL and fully-supervised baseline models are trained and evaluated on In-Distribution data from the respective dataset. The presented models are the best-performing models, based on results from the In-Distribution test data split. Once trained, they are then evaluated on out-of-distribution data originating from unseen datasets. The ‘All’ labels indicate that the PCG signals are classified based on the dataset’s original labels (P16 has only binary labels), whereas ‘Binary’ indicates that the signals are classified between normal and abnormal. The SSL-NoDs model refers to the backbone network which has not ‘seen’ data from the downstream task dataset during pretraining.
@@
While the above augmentations agree, for the most part, with the distributions and number of occurrences in Fig. 3, they also support the findings in several previous contrastive SSL works chen_simple_2020 ; gopal_3kg_2021 , in which stronger augmentations lead to improved results.
To further validate our results and to draw concrete conclusions regarding the importance of each applied transformation, we calculate Cohen’s $d$ cohen2013statistical for each implemented function. Cohen’s $d$ provides a standardized measure of the effect size of each augmentation and allows the comparison of their magnitude. Given two groups $x_{1}$ and $x_{2}$, Cohen’s $d$ is the difference in their means divided by the pooled standard deviation ($s$). Specifically, Cohen’s $d$ is defined as:
@@ 4.3 Discussion
By diving further into the results, the true benefits of the proposed method begin to show. First of all, the drop in effectiveness between in-distribution and OOD data is apparent. Additionally, in all cases the baseline models appear to have overfit their training data distribution and fail to maintain their classification ability across distinct datasets. On the other hand, the models which have been pretrained via SSL yield far superior results. In all cases, the models trained with the proposed framework surpass the baseline by 11.68% on average and even outperform it by 21.04% in the case of the PhysioNet22 OOD task. Even more so, the SSL models were surprisingly able to surpass both the effectiveness of the baseline and their ‘In-Distribution’ counterpart, in the case of the Pascal OOD evaluation. Specifically, the downstream models trained on data from PhysioNet2016 and PhysioNet2022 were able to yield better results than the fully-supervised model trained solely on Pascal data. Furthermore, even though there is a decrease in effectiveness when the OOD datasets are left out during model pretraining, the SSL models continue to outshine the baseline across the board, with the exception of the case where the downstream task is the PhysioNet2022 data and models are evaluated on OOD data from Pascal. Finally, we would like to note that the above findings are aligned with the results of our preliminary SSL method ballas_listen2yourheart_2022 submitted in the PhysioNet2022 challenge, where the proposed model was able to avoid overfitting on the available training distribution and demonstrated adequate generalization effectiveness on completely hidden data.
@@ 5 Conclusions
Acknowledgement
The work leading to these results has been funded by the European Union under Grant Agreement No. 101057821, project RELEVIUM.